OpenStudy (anonymous):

Determine the values of 'a' for which the system has no solutions, exactly one solution, or infinitely many solutions: x + 2y = 1 2x + (a^2 - 5)y = a-1

OpenStudy (anonymous):

solve \(a^2-5=4\) should do it

OpenStudy (anonymous):

how did you arrive to the = 4?

OpenStudy (anonymous):

by inspection i can tell that there is a solution when a not equal plus/minus three, and there is infinitely many solution when a=+3 and no solution when a=-2 @satellite73

OpenStudy (anonymous):

i arrived at the 4 because the first one has \(x\) and the second one has \(2x\) the first one has \(2y\) so if the second one has \(4y\) then they will be parallel

OpenStudy (anonymous):

if \(a=3\) you get \[x+2y=1\\ 2x+4y=2\] and that is the same line, so infinitely many solutions

OpenStudy (anonymous):

Thanks. I understand what you did there.

OpenStudy (anonymous):

if \(x=-3\) you get \[x+2y=1\\ 2x+4y=-4\] and that has no solution

OpenStudy (anonymous):

yw